# Scalable Message Passing Neural Networks: No Need for Attention in Large Graph Representation Learning

### v1

@@ B.1.2 Universality with the residual connection
Therefore,
[…]
Thus, the condition in equation 27 reduces further to
$\Big{(}\sum_{j=1}^{D}\,\lambda_{j}(\mathbf{W})^{2}\Big{)}^{1/2}<1.$ (29)
It only remains to understand the “typical” eigenvalues of $\mathbf{W}$ (i.e. for most realization of the random matrix $\mathbf{W}$).
Step 3 - Computing the eigenvalues of $\mathbf{W}$
First, observe that
$\Big{(}\sum_{j=1}^{D}\,\lambda_{j}(\mathbf{W})^{2}\Big{)}^{1/2}$ $\leq\Big{(}\sum_{j=1}^{D}\,\lambda_{1}(\mathbf{W})^{2}\Big{)}^{1/2}$ (30)
$=$ $\sqrt{D}|\lambda_{1}(\mathbf{W})|$ (31)
$=$ $\sqrt{D}\,s_{\operatorname{max}}(\mathbf{W})^{2}.$ (32)
Note that $W_{i,j}=\varsigma\tilde{W}_{i,j}$ where $\{\tilde{W}_{i,j}\}_{i,j=1}^{D}$ are i.i.d. standard normal random variables; meaning that the eigenvalues of $\mathbf{W}$ are exactly equal to $\varsigma$ times those of $\tilde{\mathbf{W}}$. By Gordon’s Theorem, with $N=n=D$ and $t=\sqrt{D}$, the following holds with at-least $1-e^{-D/2}$ probability:
$\operatorname{s}_{\operatorname{max}}(\tilde{\mathbf{W}})\leq 2\sqrt{D}+\sqrt{D}=3\sqrt{D}.$ (33)
Upon combining the inequalities in equation 32 and in equation 33 we deduce that: with at-least $1-e^{-D/2}$ probability the following holds
$\Big{(}\sum_{j=1}^{D}\,\lambda_{j}(\mathbf{W})^{2}\Big{)}^{1/2}=$ $\Big{(}\sum_{j=1}^{D}\,\varsigma^{2}\lambda_{j}(\tilde{\mathbf{W}})^{2}\Big{)}^{1/2}$ (34)
[…]
Note that equation 34 holds since the eigenvalues of $\mathbf{W}$ are equal to $\varsigma$ times the eigenvalues of $\tilde{\mathbf{W}}$ given that $\mathbf{W}=\varsigma\tilde{\mathbf{W}}$.
In particular, if $0<\varsigma<1/(9D^{3/2})$ then, with probability at-least $1-e^{-D/2}$ we have that
[…]
Step 4 - Conclusion
Consequentially, equation 35 implies that equation 29 is satisfied with probability at-least $1-e^{-D/2}$; meaning that, the map $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}(\mathbf{X})$ is injective with probability at-least $1-e^{-D/2}$ too. ∎
The following is a more technical version of Theorem 4.2; which shows that the result can be made to hold explicitly with high probability when appropriately sampling a weight matrix.
Theorem B.6 (Universality Regained).
Suppose that $\zeta$ satisfies Assumption B.1. Let $N,D$ be positive integers and fix a hyperparameter $0<\varsigma<1/(9D^{3/2})$. Let $\mathcal{G}$ be a complete graph on $N$ nodes, let $\tilde{\mathbf{A}}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(I_{i\sim j\vee i=j}/\sqrt{\operatorname{deg}(v_{i})\operatorname{deg}(v_{j})})_{i,j=1}^{N}$, let $\mathbf{W}$ be a $D\times D$ random matrix with i.i.d. entries $W_{i,j}\sim N(0,\varsigma^{2})$, and define $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}$ as in Lemma B.4.
Then, with probability at-least $1-e^{-D/2}$, the class of maps $\hat{f}:\mathbb{R}^{N\times D}\to\mathbb{R}$ with representation
$\hat{f}=\operatorname{MLP}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}$
is dense in $\mathcal{C}(\mathbb{R}^{N\times D},\mathbb{R})$ for the topology of uniform convergence on compacts sets, where $\operatorname{MLP}:\mathbb{R}^{N\times D}\to\mathbb{R}$ is an MLP with activation function $\zeta$.
Proof of Theorem B.6 (and thus of Theorem 4.2).
By [49, Theorem 3.1] and [42, Theorem 3.2] the class $\mathcal{MLP}_{N\times D}^{\zeta}$ is dense in $\mathcal{C}(\mathbb{R}^{N\times D},\mathbb{R})$ for the topology of uniform convergence on compacts sets since $\zeta$ satisfies Assumption B.1.
By construction, the map $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}$ is continuous. By Lemma B.4 the map is injective with probability at-least $1-e^{-D/2}$. Therefore, [48, Theorem 3.4] implies that the class
$\mathcal{MLP}_{N\times D}^{\zeta}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{\{}\operatorname{MLP}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}:\,\operatorname{MLP}\in\mathcal{MLP}_{N\times D}^{\zeta}\big{\}}$
is dense in $\mathcal{C}(\mathbb{R}^{N\times D},\mathbb{R})$ for the topology of uniform convergence on compact sets. ∎
The probability appearing in Theorem 4.2 implies that, at initialization, there is a high probability ($1-e^{-D/2}$) that the class $\mathcal{F}_{\mathcal{G},\mathbf{W}}^{\operatorname{residual}}$ is expressive enough to learn any continuous function on $\mathcal{C}(\mathbb{R}^{N\times D})$ within large domains. Note that $D$ is the hidden dimension, which tends to be high; thus, Theorem 4.2 holds with probability $\approx 1$.
Appendix C Training Details
@@
Dataset information. All datasets used in this work are publicly available. For OGB datasets, see [2]; for Pokec, see [53]; and for the spatio-temporal datasets, see PyTorch Geometric Temporal [54]. Following DIFFormer [11], we perform experiments on both image and text datasets to test the applicability of SMPNNs to a variety of tasks. The procedure described next is exactly the same as in [11]. We use STL(-10) and CIFAR(-10) as our image datasets. For the first, we use all 13,000 images, each classified into one of 10 categories. In the case of the latter, CIFAR, the dataset was originally pre-processed in previous work using 1,500 images from each of the 10 categories, resulting in a total of 15,000 images. For both datasets, 10, 50, or 100 instances per class were randomly selected for the training set (we use the same random splits as in the baselines for reproducibility and to obtain a fair comparison), 1,000 instances for validation, and the remaining instances for testing.
@@
Train, validation, and test splits. We use all public data splits when available, such as in the case of the OGB benchmarks [2]. Otherwise we follow the data split in the literature [10, 11, 12] for a consistent comparison. Note that the code for these baselines is available online, including torch seeds for datasets with random splits. We use the exact same seeds (123 in most cases).
@@
Graph Sampling and Batching for Training and Inference. For ogbn-arxiv, CIFAR, STL, 20News, Chickenpox, Covid, and Wikipath, we employ full-graph training. Specifically, we input the entire graph into the model and simultaneously predict all node labels for the loss computation using a single GPU. During inference, we also use the entire graph as input and calculate the evaluation metric based on the predictions for the validation and test set nodes. For the rest of the datasets (except ogbn-papers-100M discussed later), due to their large sizes, we adopt the mini-batch training approach of [10, 11, 12]. More concretely, the subsampling procedure of the Graph Transformer baselines uses the following approach for training: a batch of nodes from the large graph dataset is selected at random, and based on this, the induced subgraph is sampled by extracting the edges connecting the selected nodes from the full graph adjacency matrix (the induced subgraph is obtained using subgraph from torch geometric utils). For inference, the entire graph is fed into the model on CPU. For the exceptionally large ogbn-papers100M graph, which cannot fit into CPU memory, we apply the same mini-batch partition strategy used during training for inference using a neighbor sampler with 3-hops and 15, 10, and 5 neighbors per hop using the standard NeighborLoader from torch geometric, in line with [12].
@@
GPU Memory Scaling Experiments. In these experiments, we use a single Tesla V100 GPU. We use the same training configuration for all networks: node batch sizes of 10K, 20K, 40K, 60K, 80K, and 100K (using subgraph), no weight decay, Adam optimizer with a learning rate of $10^{-3}$, and dropout of $0.1$. We set all model configurations to have 256 hidden channels and 6 layers, modifying only the parts of the configuration discussed in the main text. For SGFormer, we increase the number of GNN layers to 6 and fix the linear attention layer to 1, as in the original paper [12].
@@
Model Configurations. For most of the large-scale graph experiments (unless otherwise stated), we use SMPNNs with 6 layers. Note that we test other configurations; for instance, in Table 7. For the image, text, and spatio-temporal datasets, we use 2 layers to avoid overfitting. The exact command configurations can be found in the supplementary material.

### v2

@@ Appendix C Oversmoothing and Residual Connections
In particular, the above characterization relies on analyzing message-passing from the perspective of gradient flows in the limit of infinitely many layers. Although analyzing the asymptotic behavior of GNNs is interesting from a theoretical perspective, the normalized Dirichlet energy is not directly predictive of model success, since both over-smoothing and over-sharpening will lead to degradation in performance (Giovanni et al., 2023). We provide an alternative perspective through the lens of universal approximation, which does not rely on asymptotic behavior. Specifically, we demonstrate that the universal approximation property of downstream classifiers is compromised when passing input features through graph convolution layers, but it can be restored with a residual connection.
Moreover, by the results of Riedi et al. (2023), residual connections also make the loss landscape of deep learning models less jagged, which may aid optimization (which we do not tackle in this work).
Appendix D Proofs
Proof of Theorem 4.5.
For any $\mathbf{X}\in\mathbb{R}^{N\times D}$, define its row-average $\bar{\mathbf{x}}:=\frac{1}{N}\mathbf{1}_{N}^{\top}\mathbf{X}\in\mathbb{R}^{1\times D}.$ Since $\tilde{\mathbf{A}}=J_{N}/N=(1/N)\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, we have $\tilde{\mathbf{A}}\mathbf{X}=\frac{1}{N}\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\mathbf{X}=\mathbf{1}_{N}\bar{\mathbf{x}}.$ Hence every row of $\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}$ is identical. Therefore the image of $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}$ is contained in the proper subspace $\left\{\mathbf{1}_{N}\mathbf{v}^{\top}:\;\mathbf{v}\in\mathbb{R}^{D}\right\}\subsetneq\mathbb{R}^{N\times D},$ so $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}$ is not injective. Concretely, any two inputs with the same row-average produce the same output.
By the universal approximation theorem, the class $\mathcal{MLP}_{N\times D}^{\zeta}$ is dense in $\mathcal{C}(\mathbb{R}^{N\times D})$ for the topology of uniform convergence on compact sets. The map $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}$ is continuous but not injective. Therefore, by (Kratsios and Bilokopytov, 2020, Theorem 3.4), precomposition with $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}$ does not preserve density. ∎
Proof of Lemma 4.6.
Every $\mathbf{X}\in\mathbb{R}^{N\times D}$ admits a unique decomposition $\mathbf{X}=\mathbf{1}_{N}\mathbf{u}^{\top}+\mathbf{Y}$ with $\mathbf{1}_{N}^{\top}\mathbf{Y}=0$ for some $\mathbf{u}\in\mathbb{R}^{D}$. Indeed, define the row-average vector $\mathbf{u}^{\top}=\frac{1}{N}\mathbf{1}_{N}^{\top}\mathbf{X}$ and set $\mathbf{Y}=\mathbf{X}-\mathbf{1}_{N}\mathbf{u}^{\top}$. Then $\mathbf{1}_{N}^{\top}\mathbf{Y}=0$ by construction, and the decomposition is unique because matrices of the form $\mathbf{1}_{N}\mathbf{v}^{\top}$ (which have identical rows) intersect the subspace $\{\mathbf{Y}:\mathbf{1}_{N}^{\top}\mathbf{Y}=0\}$ only at the zero matrix. Intuitively, this separates the node features into a global average signal $\mathbf{1}_{N}\mathbf{u}^{\top}$ and variations around the average $\mathbf{Y}$. Since $\tilde{\mathbf{A}}=J_{N}/N$, multiplication by $\tilde{\mathbf{A}}$ averages the rows of a matrix, yielding $\tilde{\mathbf{A}}(\mathbf{1}_{N}\mathbf{u}^{\top})=\mathbf{1}_{N}\mathbf{u}^{\top}$ and $\tilde{\mathbf{A}}\mathbf{Y}=0$. Applying the residual convolution operator $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}(\mathbf{X})=\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}+\mathbf{X}$ to the decomposition therefore gives $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}(\mathbf{X})=\mathbf{1}_{N}\mathbf{u}^{\top}\mathbf{W}+\mathbf{1}_{N}\mathbf{u}^{\top}+\mathbf{Y}=\mathbf{1}_{N}\mathbf{u}^{\top}(I_{D}+\mathbf{W})+\mathbf{Y}$. To determine injectivity it suffices to examine the kernel of this linear map. Setting $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}(\mathbf{X})=0$ yields $\mathbf{1}_{N}\mathbf{u}^{\top}(I_{D}+\mathbf{W})+\mathbf{Y}=0$. The first term has identical rows while $\mathbf{Y}$ has rows summing to zero, so they belong to complementary subspaces and must vanish separately. Hence $\mathbf{Y}=0$ and $\mathbf{u}^{\top}(I_{D}+\mathbf{W})=0$. This system has only the trivial solution $\mathbf{u}=0$ if and only if $I_{D}+\mathbf{W}$ is invertible. Since the eigenvalues of $I_{D}+\mathbf{W}$ are $1+\lambda_{i}$ where $\lambda_{i}\in\sigma(\mathbf{W})$, this occurs precisely when $1\notin\sigma(\mathbf{W})$. ∎
Proof of Theorem 4.7.
By the universal approximation theorem, $\mathcal{MLP}_{N\times D}^{\zeta}$ is dense in $\mathcal{C}(\mathbb{R}^{N\times D})$ for the topology of uniform convergence on compact sets. By Lemma 4.6, $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}$ is continuous and injective whenever $1\notin\sigma(\mathbf{W})$. Hence (Kratsios and Bilokopytov, 2020, Theorem 3.4) implies that precomposition with $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}$ preserves density.
For the probabilistic statement, the exceptional set $\{\mathbf{W}:\det(I_{D}+\mathbf{W})=0\}$ is the zero set of a nonzero polynomial in the entries of $\mathbf{W}$, so it has Lebesgue measure zero. Therefore it is hit with probability zero by any absolutely continuous distribution. ∎
Proof of Corollary 4.8.
Let $\operatorname{vec}(\cdot)$ denote the vectorization operator, which stacks the columns of a matrix $\mathbf{X}\in\mathbb{R}^{N\times D}$ into a vector in $\mathbb{R}^{ND}$. The vectorization identity $\operatorname{vec}(\mathbf{A}\mathbf{X}\mathbf{B})=(\mathbf{B}^{\top}\otimes\mathbf{A})\operatorname{vec}(\mathbf{X})$ gives $\operatorname{vec}(\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}+\mathbf{X})=(\mathbf{W}^{\top}\otimes\tilde{\mathbf{A}})\operatorname{vec}(\mathbf{X})+\operatorname{vec}(\mathbf{X})=\left(I_{ND}+\mathbf{W}^{\top}\otimes\tilde{\mathbf{A}}\right)\operatorname{vec}(\mathbf{X})$, where $\otimes$ is the Kronecker product. Thus injectivity of $\mathcal{L}^{\operatorname{conv+r}}_{\mathcal{G},\mathbf{W}}$ is equivalent to invertibility of $I_{ND}+\mathbf{W}^{\top}\otimes\tilde{\mathbf{A}}$. If $\|\tilde{\mathbf{A}}\|_{2}\,\|\mathbf{W}\|_{2}<1$, then $\|\mathbf{W}^{\top}\otimes\tilde{\mathbf{A}}\|_{2}=\|\mathbf{W}^{\top}\|_{2}\,\|\tilde{\mathbf{A}}\|_{2}=\|\mathbf{W}\|_{2}\,\|\tilde{\mathbf{A}}\|_{2}<1,$ so $I_{ND}+\mathbf{W}^{\top}\otimes\tilde{\mathbf{A}}$ is invertible by the Neumann series. Hence the feature map is injective. Since the map is linear and therefore continuous, Theorem 3.4 of Kratsios and Bilokopytov (2020) implies that the class $\mathcal{F}^{\operatorname{residual}}_{\mathcal{G},\mathbf{W}}$ is dense in $\mathcal{C}(\mathbb{R}^{N\times D})$. ∎
Appendix E Training Details
@@
Dataset information. All datasets used in this work are publicly available. For OGB datasets, see (Hu et al., 2020); for pokec, see (Leskovec and Krevl, 2014); and for the spatio-temporal datasets, see PyTorch Geometric Temporal (Rozemberczki et al., 2021). Following DIFFormer (Wu et al., 2023a), we perform experiments on both image and text datasets to test the applicability of SMPNNs to a variety of tasks. The procedure described next is exactly the same as in (Wu et al., 2023a). We use STL(-10) and CIFAR(-10) as our image datasets. For the first, we use all 13,000 images, each classified into one of 10 categories. In the case of the latter, CIFAR, the dataset was originally pre-processed in previous work using 1,500 images from each of the 10 categories, resulting in a total of 15,000 images. For both datasets, 10, 50, or 100 instances per class were randomly selected for the training set (we use the same random splits as in the baselines for reproducibility and to obtain a fair comparison), 1,000 instances for validation, and the remaining instances for testing.
@@
Train, validation, and test splits. We use all public data splits when available, such as in the case of the OGB benchmarks (Hu et al., 2020). Otherwise we follow the data split in the literature (Wu et al., 2022; 2023a; 2023b) for a consistent comparison. Note that the code for these baselines is available online, including torch seeds for datasets with random splits. We use the exact same seeds (123 in most cases).
@@
Graph Sampling and Batching for Training and Inference. For ogbn-arxiv, CIFAR, STL, 20News, Chickenpox, Covid, and WikiMath, we employ full-graph training. Specifically, we input the entire graph into the model and simultaneously predict all node labels for the loss computation using a single GPU. During inference, we also use the entire graph as input and calculate the evaluation metric based on the predictions for the validation and test set nodes. For the rest of the datasets (except ogbn-papers-100M discussed later), due to their large sizes, we adopt the mini-batch training approach of (Wu et al., 2022; 2023a; 2023b). More concretely, the subsampling procedure of the Graph Transformer baselines uses the following approach for training: a batch of nodes from the large graph dataset is selected at random, and based on this, the induced subgraph is sampled by extracting the edges connecting the selected nodes from the full graph adjacency matrix (the induced subgraph is obtained using subgraph from torch geometric utils). For inference, the entire graph is fed into the model on CPU. For the exceptionally large ogbn-papers-100M graph, which cannot fit into CPU memory, we apply the same mini-batch partition strategy used during training for inference using a neighbor sampler with 3-hops and 15, 10, and 5 neighbors per hop using the standard NeighborLoader from torch geometric, in line with (Wu et al., 2023b).
@@
GPU Memory Scaling Experiments. In these experiments, we use a single Tesla V100 GPU. We use the same training configuration for all networks: node batch sizes of 10K, 20K, 40K, 60K, 80K, and 100K (using subgraph), no weight decay, Adam optimizer with a learning rate of $10^{-3}$, and dropout of $0.1$. We set all model configurations to have 256 hidden channels and 6 layers, modifying only the parts of the configuration discussed in the main text. For SGFormer, we increase the number of GNN layers to 6 and fix the linear attention layer to 1, as in the original paper (Wu et al., 2023b).
@@
Model Configurations. For most of the large-scale graph experiments (unless otherwise stated), we use SMPNNs with 6 layers. Note that we test other configurations; for instance, in Table 7. For the image, text, and spatio-temporal datasets, we use 2 layers to avoid overfitting. The exact configurations can be found in the supplementary material.
Appendix F GPU Memory Scaling Plots